# label=•, leftmargin=2em, nolistsep

### v1

@@ Abstract
There is a significant relevance of federated learning (FL) in the realm of Artificial Intelligence of Things (AIoT). However, most of existing FL works are not conducted on datasets collected from authentic IoT devices that capture unique modalities and inherent challenges of IoT data. In this work, we introduce FedAIoT, an FL benchmark for AIoT to fill this critical gap. FedAIoT includes eight datatsets collected from a wide range of IoT devices. These datasets cover unique IoT modalities and target representative applications of AIoT. FedAIoT also includes a unified end-to-end FL framework for AIoT that simplifies benchmarking the performance of the datasets. Our benchmark results shed light on the opportunities and challenges of FL for AIoT. We hope FedAIoT could serve as an invaluable resource to foster advancements in the important field of FL for AIoT. The repository of FedAIoT is maintained at https://github.com/AIoT-MLSys-Lab/FedAIoT.
@@ 1 Introduction
The proliferation of the Internet of Things (IoT) such as smartphones, drones, and sensors deployed at homes, as well as the gigantic amount of data they capture, have revolutionized the way we work, live, and interact with the world. The advances in Artificial Intelligence (AI) have boosted the integration of IoT and AI that turns Artificial Intelligence of Things (AIoT) into reality. However, data captured by IoT devices usually contain privacy-sensitive information. In recent years, federated learning (FL) has emerged as a privacy-preserving solution that enables extracting knowledge from the collected data while keeping the data on the devices (Kairouz et al., 2021; Wang et al., 2021).
Despite the significant relevance of FL in the realm of AIoT, most existing FL works are conducted on well-known datasets such as CIFAR-10 and CIFAR-100. These datasets, however, do not originate from authentic IoT devices and thus fail to capture the unique modalities and inherent challenges associated with real-world IoT data. This notable discrepancy underscores a strong need for an IoT-oriented FL benchmark to fill this critical gap.
In this work, we present FedAIoT, an FL benchmark for AIoT. At its core, FedAIoT includes eight well-chosen datasets collected from a wide range of authentic IoT devices from smartwatch, smartphone and Wi-Fi routers, to drones, smart home sensors, and head-mounted device that either have already become an indispensable part of people’s daily lives or are driving emerging applications. These datasets encapsulate a variety of unique IoT-specific data modalities such as wireless data, drone images, and smart home sensor data (e.g., motion, energy, humidity, temperature) that have not been explored in existing FL benchmarks. Moreover, these datasets target some of the most representative applications and innovative use cases of AIoT that are not possible with other technologies.
@@
We have conducted a systematic benchmarking on the eight datasets using the end-to-end framework. Specifically, we examine the impact of varying degrees of non-IID data distributions, FL optimizers, and client sampling ratios on the performance of FL. We also evaluate the impact of noisy labels, a prevalent challenge in IoT datasets, as well as the effects of quantized training, a technique that tackles the practical limitation of resource-constrained IoT devices. Our benchmark results provide valuable insights into both the opportunities and challenges of FL for AIoT. Given the significant relevance of FL in the realm of AIoT, we hope FedAIoT could act as a valuable tool to foster advancements in the important area of FL for AIoT.
In summary, our work makes the following contributions.
@@ 2 Related Work
The importance of data to FL research pushes the development of FL benchmarks on a variety of data modalities. Existing FL benchmarks, however, predominantly center around curating FL datasets in the domain of computer vision (CV) (He et al., 2021b; Dimitriadis et al., 2022), natural language processing (NLP) Lin et al. (2021); Dimitriadis et al. (2022), medical imaging (Terrail et al., 2022), speech and audio (Zhang et al., 2023a; Dimitriadis et al., 2022), and graph neural networks (He et al., 2021a). For example, FedCV He et al. (2021b), FedNLP (Lin et al., 2021), and FedAudio (Zhang et al., 2023a) focuses on benchmarking CV, NLP, and audio-related datasets and tasks respectively; FLUTE (Dimitriadis et al., 2022) covers a mix of datasets from CV, NLP, and audio; FLamby (Terrail et al., 2022) primarily focuses on medical imaging datasets; and FedMultimodal (Feng et al., 2023) includes multimodal datasets in the domain of emotion recognition, healthcare, multimedia, and social media. As summarized in Table 1, although these benchmarks have significantly contributed to FL research, a dedicated FL benchmark explicitly tailored for IoT data is absent. Compared to these existing FL benchmarks, FedAIoT is specifically designed to fill this critical gap by providing a dedicated FL benchmark on data collected from a wide range of authentic IoT devices.
Table 1: Comparison between FedAIoT with existing FL benchmarks.
Data Type FL Framework Designed for IoT Noisy Labels Quantized Training
FLamby Medical Images
No
No No
FedAudio Audio Data
No
Uniform No
FedMultimodal Multimodality Data
No
Uniform No
FedGraphNN Graph Data
No
No No
FedNLP Natural Language
No
No No
FLUTE Images and Text
No
No Server Side only
FedCV Images
No
No No
FedAIoT IoT Data Yes Probabilistic Server and Client Side
@@ 3 Design of FedAIoT
3.1 Datasets
Objective: The objective of FedAIoT is to provide a benchmark that consists of high-quality and well-validated datasets collected from a wide range of IoT devices, sensor modalities, and applications. We have made significant efforts to validate a much larger pool of existing datasets and identified the high-quality ones to include in our benchmark. Table 2 provides an overview of the eight datasets included in FedAIoT.These datasets have diverse sizes (small, medium, large). The rationale behind this design choice is to accommodate researchers who have different computing resources. For example, researchers who have very limited computing resources can still pick the relatively small datasets included in our benchmark to develop and evaluate their algorithms. In this section, we provide a brief overview of each included dataset.
WISDM: The Wireless Sensor Data Mining (WISDM) dataset (Weiss et al., 2019; Lockhart et al., 2011) is one of the widely used datasets for the task of daily activity recognition using accelerometer and gyroscope sensor data collected from smartphones and smartwatches. WISDM includes data collected from $51$ participants performing $18$ daily activities, each in a $3$-minute session. We combined activities such as eating soup, chips, pasta, and sandwiches into a single category called “eating”, and removed uncommon activities related to playing with balls, such as kicking, catching, or dribbling. We randomly selected $45$ participants as the training set and the rest of the $6$ participants were assigned to the test set. Given that in real-life settings individuals may not carry a smartphone and wear a smartwatch at the same time, we partition WISDM into two separate datasets: WISDM-W with smartwatch data only and WISDM-P with smartphone data only. The total number of samples in the training and test set is $16,569$ and $4,103$ for WISDM-W and $13,714$ and $4,073$ for WISDM-P respectively. No Licence was explicitly mentioned on the dataset homepage.
Widar: The Widar dataset (Yang, 2020; Zheng et al., 2019) is designed for contactless gesture recognition using Wi-Fi signal strength measurements collected from strategically placed access points. This system uses an Intel $5300$ NIC with $3\times 3$ antenna pairs. The dataset features data from $17$ participants showcasing $22$ unique gestures, including actions like push, pull, sweeping, and clapping. To maintain consistency between training and test sets, only gestures recorded by more than three users are retained. This decision ensures the test set doesn’t have data from the same participants as the training set. As a result, the balanced dataset includes nine gestures with $11,372$ samples in the training set and $5,222$ in the test set. The dataset is licensed under the Creative Commons Attribution-NonCommercial 4.0 International Licence (CC BY 4).
CASAS: The CASAS dataset (Schmitter-Edgecombe & Cook, 2009), derived from the CASAS smart home project, is a smart home sensor dataset for the task of recognizing activities of daily living (ADL) based on sequences of sensor states over time to support the application of independent living. Data were collected from three distinct apartments, each equipped with three types of sensors: motion sensors, temperature sensors, and door sensors. We have selected five specific datasets from CASAS named “Milan”, “Cairo”, “Kyoto2”, “Kyoto3”, and “Kyoto4” based on the uniformity of their sensor data representation. The original ADL categories within each dataset have been consolidated into $11$ categories related to home activities such as “sleep”, “eat”, and “bath”. Activities not fitting within these categories were collectively classified as “other”. The training and test set was made using an 80-20 split. Each data sample is a categorical time series of length $2,000$, representing sensor states over a certain period of time. The total number of samples in the training and test set is $12,190$ and $3,048$ respectively. No Licence was explicitly mentioned on the dataset homepage.
@@ 3.1 Datasets
Dataset IoT Platform Data Modality Data Dimension Dataset Size # Training Samples # Clients
@@
UT-HAR Wi-Fi Router Wireless Signal $3\times 30\times 250$ 854 MB $3,977$
20
Widar Wi-Fi Router Wireless Signal $22\times 20\times 20$ 3.3 GB $11,372$
40
VisDrone Drone Images $3\times 224\times 224$ 1.8 GB $6,471$
30
@@
EPIC-SOUNDS Augmented Reality Acoustics $400\times 128$ 34 GB $60,055$
210
AEP: The Appliances Energy Prediction (AEP) dataset (Candanedo et al., 2017) is another smart home sensor dataset but designed for the task of home energy usage prediction. Data were collected from energy sensors, temperature sensors, and humidity sensors installed inside a home every $10$ minutes over $4.5$ months. The total number of samples in the training and test set is $15,788$ and $3,947$ respectively. No Licence was explicitly mentioned on the dataset homepage.
@@ 3.2.1 Non-IID Data Partitioning
A key characteristic of FL is that data distribution at different clients is non-IID. The objective of non-IID data partitioning is to partition the training set such that data allocated to different clients follow the non-IID distribution. The eight datasets included in FedAIoT cover three fundamental tasks: classification, regression, and object detection. FedAIoT incorporates three different non-IID data partitioning schemes that are designed for the three tasks respectively.
@@
Scheme#1: Non-IID Partition over Output Labels. For the task of classification (WISDM-W, WISDM-P, UT-HAR, Widar, CASAS, EPIC-SOUNDS) with $C$ classes, we first generate a distribution over the classes for each client by drawing from a Dirichlet distribution with parameter $\alpha$ (Hsu et al., 2019), where lower values of $\alpha$ generate more skewed distribution favoring a few classes whereas higher values of $\alpha$ result in more balanced class distributions. We use the same $\alpha$ to determine the number of samples each client receives. In addition, by drawing from a Dirichlet distribution with parameter $\alpha$, we create a distribution over the total number of samples, which is then used to allocate a varying number of samples to each client, where lower values of $\alpha$ lead to a few clients holding a majority of the samples whereas higher values of $\alpha$ create a more balanced distribution of samples across clients. Therefore, this approach allows us to generate non-IID data partitions both the class distribution and the number of samples can vary across the clients.
@@
Scheme#2: Non-IID Partition over Input Features. The task of object detection (VisDrone) does not have specific classes. In such case, we use the input features to create non-IID partitions. Specifically, similar to He et al. (2021b), we first used ImageNet (Russakovsky et al., 2015) features generated from the VGG19 model (Liu & Deng, 2015), which encapsulate visual information required for subsequent analysis. With these ImageNet features as inputs, we performed clustering in the feature space using $K$-nearest neighbors to partition the dataset into clusters. Each cluster is a pseudo-class, representing a set of images sharing common visual characteristics as per the extracted ImageNet features. Lastly, Dirichlet allocation was applied on top of the pseudo-classes to create the non-IID distribution across different clients.
@@
Scheme#3: Non-IID Partition over Output Distribution. For the task of regression (AEP) where output is characterized as a continuous variable, we utilize Quantile Binning (Pyle, 1999). Specifically, we divide the range of the output variable into $10$ equal groups or quantiles, ensuring that each bin accommodates roughly the same number of samples. Each category or bin is treated as a pseudo-class. Lastly, we apply Dirichlet allocation to generate the non-IID distribution of data across the clients.
@@ 3.2.2 IoT-specific Data Preprocessing
WISDM: We followed the standard preprocessing techniques used in accelerometer and gyroscope-based activity recognition for WISDM Ravi et al. (2005); Reyes-Ortiz et al. (2016); Ronao & Cho (2016). Specifically, for each $3$-minute session, we used a $10$-second sliding window with $50$% overlap to extract samples from the raw accelerometer and gyroscope data sequences. We then normalize each dimension of the extracted samples by removing the mean and scaling to unit variance.
@@ 3.2.3 IoT-friendly Models
Since FedAIoT focuses on IoT devices which are resource-constrained, our choice of models is informed by a combination of model accuracies (Reinbothe, 2023; Yang et al., 2023; Terven & Cordova-Esparza, 2023; Liciotti et al., 2019; Seyedzadeh et al., 2018; Sholahudin et al., 2016; Huh et al., 2023) and model efficiency. For each included dataset, we evaluated multiple model candidates, and selected the best-performing one that adheres to the resource constraint of each IoT device listed in Table 8. As an example, for UT-HAR dataset, two model candidates (ViT and ResNet18) have similar accuracy, and we chose the more efficient ResNet18 in our benchmark. Table 3 lists the selected model for each dataset. The detail of the architecture of each model is described in Appendix A.3.
Dataset WISDM-W WISDM-P UT-HAR Widar VisDrone CASAS AEP EPIC-SOUNDS
Partition Output Output Output Output Input Output Output Output
Labels Labels Labels Labels Features Labels Distribution Labels
Model LSTM LSTM ResNet18 ResNet18 YOLOv8n BiLSTM MLP ResNet18

### v2

# FedAIoT: A Federated Learning Benchmark for Artificial Intelligence of Things

@@ Abstract
There is a significant relevance of federated learning (FL) in the realm of Artificial Intelligence of Things (AIoT). However, most existing FL works do not use datasets collected from authentic IoT devices and thus do not capture unique modalities and inherent challenges of IoT data. To fill this critical gap, in this work, we introduce FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of IoT devices. These datasets cover unique IoT modalities and target representative applications of AIoT. FedAIoT also includes a unified end-to-end FL framework for AIoT that simplifies benchmarking the performance of the datasets. Our benchmark results shed light on the opportunities and challenges of FL for AIoT. We hope FedAIoT could serve as an invaluable resource to foster advancements in the important field of FL for AIoT. The repository of FedAIoT is maintained at https://github.com/AIoT-MLSys-Lab/FedAIoT.
Keywords: Benchmarks, Federated Learning, Artificial Intelligence of Things, AIoT
@@ 1 Introduction
The proliferation of the Internet of Things (IoT) such as smartphones, drones, and sensors deployed at homes, and the gigantic amount of data they capture, have revolutionized the way we work, live, and interact with the world. The advances in Artificial Intelligence (AI) have boosted the integration of IoT and AI that turns Artificial Intelligence of Things (AIoT) into reality. However, data captured by IoT devices usually contain privacy-sensitive data. In recent years, federated learning (FL) has emerged as a privacy-preserving solution that enables extracting knowledge from the collected data while keeping the data on the devices (Kairouz et al., 2021; Wang et al., 2021).
Despite the significant relevance of FL, most existing FL works are conducted on well-known datasets. We present an overview of the datasets different FL papers have used from the pioneering works to more recent publications in Table 9 These datasets, however, do not originate from authentic IoT devices and thus fail to capture the unique modalities and inherent challenges associated with real-world IoT data. This discrepancy underscores a strong need for an IoT-oriented FL benchmark to fill this critical gap.
In this work, we present FedAIoT, an FL benchmark for AIoT (Figure 1). At its core, FedAIoT includes eight well-chosen datasets collected from a wide range of IoT devices from smartwatch, smartphone and Wi-Fi routers, to drones, smart home sensors, and head-mounted device that either have already become an indispensable part of people’s daily lives or are driving emerging applications. These datasets encapsulate a variety of unique IoT-specific data modalities such as wireless data, drone images, and smart home sensor data (e.g., motion, energy, humidity, temperature) that have not been explored in existing FL benchmarks. Moreover, these datasets target some of the most representative applications and innovative use cases of AIoT that are not possible with other technologies.
@@
We have conducted a systematic benchmarking on the eight datasets using the end-to-end framework. Specifically, we examine the impact of varying degrees of non-IID data distributions, FL optimizers, and client sampling ratios on the performance of FL. We also evaluate the impact of noisy labels, a prevalent challenge in IoT datasets, as well as the effects of quantized training, a technique that tackles the practical limitation of resource-constrained IoT devices. Our benchmark results provide valuable insights into both the opportunities and challenges of FL for AIoT. Given the significant relevance of FL in the realm of AIoT, we hope FedAIoT could act as a valuable tool to foster advancements in the important area of FL for AIoT. In summary, our work makes the following contributions.
Table 1: Comparison between FedAIoT with existing FL benchmarks.
{tblr}
columneven = c, column3 = c, column5 = c, hline1,10 = -0.08em, hline2 = -, & Data Type FL Framework Designed for IoT Noisy LabelS Quantized Training
FLamby Medical Images No No No
FedAudio Audio Data No Uniform  No
FedMultimodal Multimodality Data No Uniform No
FedGraphNN Graph Data No No No
FedNLP Natural Language No No No
FLUTE Images and Text No No Server Side only
FedCV Images No No No
FedAIoT IoT Data Yes Probabilistic Server and Client Side
@@ 2 Related Work
The importance of data to FL research pushes the development of FL benchmarks on a variety of data modalities. Existing FL benchmarks, however, predominantly center around curating FL datasets in the domain of computer vision (CV) (He et al., 2021b; Dimitriadis et al., 2022), natural language processing (NLP) Lin et al. (2021); Dimitriadis et al. (2022), medical imaging (Terrail et al., 2022), speech and audio (Zhang et al., 2023a; Dimitriadis et al., 2022), and graph neural networks (He et al., 2021a). For example, FedCV He et al. (2021b), FedNLP (Lin et al., 2021), and FedAudio (Zhang et al., 2023a) focuses on benchmarking CV, NLP, and audio-related datasets and tasks respectively; FLUTE (Dimitriadis et al., 2022) covers a mix of datasets from CV, NLP, and audio; FLamby (Terrail et al., 2022) primarily focuses on medical imaging datasets; and FedMultimodal (Feng et al., 2023) includes multimodal datasets in the domain of emotion recognition, healthcare, multimedia, and social media. As summarized in Table 1, although these benchmarks have contributed to FL research, a dedicated FL benchmark explicitly tailored for IoT data is absent. Compared to these existing FL benchmarks, FedAIoT is specifically designed to fill this critical gap by providing a dedicated FL benchmark on data collected from a wide range of authentic IoT devices.
@@ 3 Design of FedAIoT
{tblr}
cells = c, column3 = t, column4 = t, hline1-2,10 = -, Dataset & IoT Platform Data Modality Data Dimension Dataset Size # Training Samples # Clients
@@
UT-HAR Wi-Fi Router Wireless Signal $3\times 30\times 250$ 854 MB $3,977$ 20
Widar Wi-Fi Router Wireless Signal $22\times 20\times 20$ 3.3 GB $11,372$ 40
VisDrone Drone Images $3\times 224\times 224$ 1.8 GB $6,471$ 30
@@
EPIC-SOUNDS Augmented Reality Acoustics $400\times 128$ 34 GB $60,055$ 210
3.1 Datasets
Objective: The objective of FedAIoT is to provide a benchmark that consists of high-quality and well-validated datasets collected from a wide range of IoT devices, sensor modalities, and applications. We have made significant efforts to validate a much larger pool of existing datasets and identified the high-quality ones to include in our benchmark. Table 2 provides an overview of the eight datasets included in FedAIoT. These datasets have diverse sizes (small: less than 7k samples; medium: 11k to 16k samples; and large: more than 60k samples). The rationale behind this design choice is to accommodate researchers who have different computing resources. For example, researchers who have very limited computing resources can still pick the relatively small datasets included in our benchmark to develop and evaluate their algorithms. In this section, we provide a brief overview of each included dataset.
WISDM: The Wireless Sensor Data Mining (WISDM) dataset (Weiss et al., 2019; Lockhart et al., 2011) is one of the widely used datasets for the task of daily activity recognition using accelerometer and gyroscope sensor data collected from smartphones and smartwatches. WISDM includes data collected from $51$ participants performing $18$ daily activities, each in a $3$-minute session. We combined activities such as eating soup, chips, pasta, and sandwiches into a single category called “eating”, and removed uncommon activities related to playing with balls, such as kicking, catching, or dribbling. We randomly selected $45$ participants as the training set and the remaining six participants were assigned to the test set. Given that the smartwatch data and smartphone data were not collected simultaneously for most subjects and thus were not synchronized precisely, we partition WISDM into two independent datasets: WISDM-W with smartwatch data only and WISDM-P with smartphone data only. The total number of samples in the training and test set is $16,569$ and $4,103$ for WISDM-W and $13,714$ and $4,073$ for WISDM-P respectively. No Licence was explicitly mentioned on the dataset homepage.
Widar: The Widar dataset (Yang, 2020; Zheng et al., 2019) is designed for contactless gesture recognition using Wi-Fi signal strength measurements collected from strategically placed access points. This system uses an Intel $5300$ NIC with $3\times 3$ antenna pairs. The dataset features data from $17$ participants showcasing $22$ unique gestures, including actions like push, pull, sweeping, and clapping. To maintain consistency between training and test sets, only gestures recorded from more than three users are retained. This decision ensures the test set doesn’t have data from the same participants as the training set. As a result, the balanced dataset includes nine gestures with $11,372$ samples in the training set and $5,222$ in the test set. The dataset is licensed under the Creative Commons Attribution-NonCommercial 4.0 International Licence (CC BY 4).
CASAS: The CASAS dataset (Schmitter-Edgecombe and Cook, 2009), derived from the CASAS smart home project, is a smart home sensor dataset for the task of recognizing activities of daily living (ADL) based on sequences of sensor states over time to support the application of independent living. Data were collected from three distinct apartments, each equipped with three types of sensors: motion sensors, temperature sensors, and door sensors. We have selected five specific datasets from CASAS named “Milan”, “Cairo”, “Kyoto2”, “Kyoto3”, and “Kyoto4” based on the uniformity of their sensor data representation. The original ADL categories within each dataset have been consolidated into $11$ categories related to home activities such as “sleep”, “eat”, and “bath”. Activities not fitting within these categories were collectively classified as “other”. The training and test set was made using an 80-20 split. Each data sample is a categorical time series of length $2,000$, representing sensor states over a certain period of time. The total number of samples in the training and test set is $12,190$ and $3,048$ respectively. No Licence was mentioned on the dataset homepage.
AEP: The Appliances Energy Prediction (AEP) dataset (Candanedo et al., 2017) is another smart home sensor dataset but designed for the task of home energy usage prediction. Data were collected from energy sensors, temperature sensors, and humidity sensors installed inside a home every $10$ minutes over $4.5$ months. The number of samples in the training and test set is $15,788$ and $3,947$ respectively. No Licence was mentioned on the dataset homepage.
@@ 3.2.1 Non-IID Data Partitioning
A key characteristic of FL is that data distribution at different clients is non-IID. The objective of non-IID data partitioning is to partition the training set such that data allocated to different clients follow the non-IID distribution. The eight datasets included in FedAIoT cover three fundamental tasks: classification, regression, and object detection. As summarized in Table 3, FedAIoT incorporates three different non-IID data partitioning schemes that are designed for the three tasks respectively.
{tblr}
cells = c, cell21 = r=2, hline1,5 = -0.08em, hline2 = -, Dataset & WISDM-W WISDM-P UT-HAR Widar VisDrone CASAS AEP EPIC-SOUNDS
Partition
Output Output Output Output Input Output Output Output
Labels Labels Labels Labels Features Labels Distribution Labels
Model LSTM LSTM ResNet18 ResNet18 YOLOv8n BiLSTM MLP ResNet18
@@
Scheme#1: Non-IID Partition over Output Labels. For the task of classification (WISDM-W, WISDM-P, UT-HAR, Widar, CASAS, EPIC-SOUNDS) with $C$ classes, we first generate a distribution over the classes for each client by drawing from a Dirichlet distribution (formal definition in Appendix A.2) with parameter $\alpha$ (Hsu et al., 2019), where lower values of $\alpha$ generate more skewed distribution favoring a few classes whereas higher values of $\alpha$ result in more balanced class distributions. We use the same $\alpha$ to determine the number of samples each client receives. In addition, by drawing from a Dirichlet distribution with parameter $\alpha$, we create a distribution over the total number of samples, which is then used to allocate a varying number of samples to each client, where lower values of $\alpha$ lead to a few clients holding a majority of the samples whereas higher values of $\alpha$ create a more balanced distribution of samples across clients. Therefore, this approach allows us to generate non-IID data partitions such that both class distribution and the number of samples can vary across the clients.
@@
Scheme#2: Non-IID Partition over Input Features. The task of object detection (VisDrone) does not have specific classes. In such case, we use the input features to create non-IID partitions. Specifically, similar to He et al. (2021b), we first used ImageNet (Russakovsky et al., 2015) features generated from the VGG19 model (Liu and Deng, 2015), which encapsulate visual information required for subsequent analysis. With these ImageNet features as inputs, we performed clustering in the feature space using $K$-nearest neighbors to partition the dataset into clusters. Each cluster is a pseudo-class, representing a set of images sharing common visual characteristics as per the extracted ImageNet features. Lastly, Dirichlet allocation was applied on top of the pseudo-classes to create the non-IID distribution across different clients.
@@
Scheme#3: Non-IID Partition over Output Distribution. For the task of regression (AEP) where output is characterized as a continuous variable, we utilize Quantile Binning (Pyle, 1999). Specifically, we divide the range of the output variable into equal groups or quantiles, ensuring that each bin accommodates roughly the same number of samples. Each category or bin is treated as a pseudo-class. Note that the number of quantiles can be set to any value in our framework. We used ten as an example to demonstrate the results. Lastly, we apply Dirichlet allocation to generate the non-IID distribution of data across the clients.
@@ 3.2.2 IoT-specific Data Preprocessing
WISDM: We followed the standard preprocessing techniques used in accelerometer and gyroscope-based activity recognition for WISDM Ravi et al. (2005); Reyes-Ortiz et al. (2016); Ronao and Cho (2016). Specifically, for each $3$-minute session, we used a $10$-second sliding window with $50$% overlap to extract samples from the raw accelerometer and gyroscope data sequences. We then normalize each dimension of the extracted samples by removing the mean and scaling to unit variance.
@@ 3.2.3 IoT-friendly Models
Since FedAIoT focuses on IoT devices which are resource-constrained, our choice of models is informed by a combination of model accuracies (Reinbothe, 2023; Yang et al., 2023; Terven and Cordova-Esparza, 2023; Liciotti et al., 2019; Seyedzadeh et al., 2018; Sholahudin et al., 2016; Huh et al., 2023) and model efficiency. For each included dataset, we evaluated multiple model candidates, and selected the best-performing ones that adhere to the resource constraints of representative IoT devices listed in Table 4. As an example, for UT-HAR dataset, two model candidates (ViT and ResNet18) have similar accuracy, and we chose the more efficient ResNet18 in our benchmark. Table 3 lists the selected model for each dataset. The detail of the architecture of each model is described in Appendix A.5.
Table 4: Representative IoT devices for each dataset.
{tblr}
cells = c, cell22 = r=3, hline1-2,5-10 = -, Dataset & Application IoT Platform Representative Devices Hardware RAM Size
WISDM-W Activity Recognition Smartwatch Apple Watch 8 512 MB to 1 GB
WISDM-P Smartphone iPhone 14 6 GB
UT-HAR Wi-Fi Router TP-Link AX1800 64 MB to 1 GB
Widar Gesture Recognition Wi-Fi Router TP-Link AX1800 64 MB to 1 GB
CASAS Independent Living Smart Home Raspberry Pi 4 1 GB to 8 GB
AEP Energy Prediction Smart Home Raspberry Pi 4 1 GB to 8 GB
VisDrone Objective Detection Drone Dji Mavic 3 + Raspberry Pi 4 1 GB to 8 GB
EPIC-SOUNDS Augmented Reality Head-mounted Device GoPro / AR Headset 1 GB to 8 GB